# Benefits of Mutual Coupling in Dynamic Metasurface Antennas for Optimizing Wireless Communications – Theory and Experimental Validation

### v1

@@ Abstract
Dynamic metasurface antennas (DMAs) are a promising embodiment of next-generation reconfigurable antenna technology to realize base stations and access points with reduced cost and power consumption. A DMA is a thin structure patterned on its front with reconfigurable radiating metamaterial elements (meta-atoms) that are excited by waveguides or cavities. Mutual coupling between the meta-atoms can result in a strongly non-linear dependence of the DMA’s radiation pattern on the configuration of its meta-atoms. However, besides the obvious algorithmic challenges of working with physics-compliant DMA models, it remains unclear how mutual coupling in DMAs influences the ability to achieve a desired wireless functionality. In this paper, we provide theoretical, numerical and experimental evidence that strong mutual coupling in DMAs increases the radiation pattern sensitivity to the DMA configuration and thereby boosts the available control over the radiation pattern, improving the ability to tailor the radiation pattern to the requirements of a desired wireless functionality. Counterintuitively, we hence encourage next-generation DMA implementations to enhance (rather than suppress) mutual coupling, in combination with suitable physics-compliant modeling and optimization. We expect the unveiled mechanism by which mutual coupling boosts the radiation pattern control to also apply to other reconfigurable antenna systems based on tunable lumped elements.
Index Terms:
Dynamic metasurface antenna, fluid antenna system, mutual coupling, physics-compliant model, coupled-dipole formalism, tunable lumped elements, optimization, full-wave simulation, 20 GHz prototype.
@@ I Introduction
However, the scalability of such multi-antenna systems is fundamentally limited by considerations of cost, power consumption, weight and footprint. To overcome these limitations, next-generation reconfigurable antenna technology seeks to exploit additional degrees of freedom in the physical layer. A fluid antenna system (FAS) has software-controllable electromagnetic properties that enable the dynamic reconfiguration of its radiation pattern [3, 4, 5, 6, 7]. Embodiments of the FAS concept range from implementations based on liquid metals offering continuous mechanical tunability to implementations based on tunable lumped elements offering electrical tunability at spatially discrete locations [5]. The realm of FAS based on tunable lumped elements is itself vast and can be classified according to whether the tunable lumped elements are themselves the primary mechanism by which the FAS radiates. For instance, reconfigurable pixel-based antennas are grids of metallic patches interconnected by tunable lumped elements [8, 9, 10, 11, 12, 13], where the patches are the main radiators and the tunable lumped elements control their coupling. In contrast, dynamic metasurface antennas (DMAs) are surfaces patterned with radiating reconfigurable meta-atoms that are excited by waveguides or cavities [14, 15, 16, 17, 18, 19, 20]; in DMAs, the tunable meta-atoms are simultaneously the tuning mechanism and the radiation mechanism. We focus for concreteness on DMAs in this paper; however, we expect the fundamental principles that we unveil to apply more broadly to any FAS based on tunable lumped elements.
@@
It is relatively well understood that mutual coupling between the reconfigurable meta-atoms [^1] The question of how mutual coupling between physically tunable elements impacts the influence of these physical-layer degrees-of-freedom on system performance is clearly different from the well-studied impact of mutual coupling in static arrays on the performance of systems for which only the coherent excitation signals are adjustable [21, 22]. profoundly influences the mathematical nature of the mapping from DMA configuration to DMA radiation pattern. Specifically, this mapping is generally non-linear due to mutual coupling, as seen in physics-compliant models for various DMA hardware designs [23, 24, 25, 26]. Because the mapping’s non-linearity severely complicates the algorithmic design, mutual coupling is generally considered a vexing nuance. Wireless practitioners hence either avoid these complications by neglecting or deliberately minimizing mutual coupling in DMAs to achieve an effectively affine mapping [27, 15, 28, 29, 19, 18, 30, 31, 20, 32], or they develop advanced algorithms based on physics-compliant models that capture mutual coupling in DMAs [17, 33, 34]. However, the influence of mutual coupling on the strength of the dependence of the DMA radiation pattern on the DMA configuration is generally not considered.
Recently, inspired by earlier work on the benefits of multi-path propagation in wireless localization [35], we presented preliminary full-wave numerical evidence that mutual coupling boosts the influence of the DMA configuration on the DMA radiation pattern [36]; we further suggested that stronger mutual coupling should therefore improve the ability to tailor the DMA radiation pattern to a desired wireless functionality [36]. This generally overlooked effect of mutual coupling in DMAs can drastically change the assessment of the role of mutual coupling in DMAs. In essence, a trade-off between strength and simplicity of the dependence of the DMA radiation pattern on the DMA configuration emerges. Intuitively, these two effects of mutual coupling can be understood in terms of multi-bounce paths between the meta-atoms [^2] The relation between physics-compliant models and a multi-bounce-path picture has been worked out in the context of RIS-parametrized radio environments using a coupled-dipole formalism in [37] and using multi-port network theory in Sec. II.A of [38].: The stronger the mutual coupling in the DMA is, the more the wave bounces back and forth between meta-atoms, resulting in an accumulation of more sensitivity to the DMA configuration, which in turn yields a i) stronger and ii) more strongly non-linear (i.e., less simple) dependence of the DMA radiation pattern on the DMA configuration. Consequently, it may be desirable to deploy DMA hardware with enhanced (rather than suppressed) mutual coupling in combination with suitable physics-compliant models and optimization techniques.
@@
    mutual coupling strength between the DMA’s meta-atoms,
@@
  • We formulate a compact, generic (implementation-agnostic), physics-compliant DMA model. We examine the mathematical role of mutual coupling therein, and clarify how to recover common simplified linear models upon neglecting mutual coupling.
@@
  • We derive analytical expressions and upper bounds for the normalized Jacobian and Hessian of the DMA radiation pattern with respect to the DMA configuration. The upper bounds monotonically increase with the mutual coupling strength.
@@
  • We systematically examine in full-wave simulations how the four pivotal quantities listed above depend on the mutual coupling strength in a DMA. We reveal clear trends of positive correlation between mutual coupling strength and sensitivity, achieved channel gain enhancement, and non-linearity of the mapping from configuration to radiation pattern.
@@
  • We conduct experiments with a DMA prototype operating in the lower K-band with 1-bit programmable meta-atoms. We introduce an original time-gating technique to adjust the effective mutual coupling strength in post-processing. We report clear experimental evidence for the numerically observed trends.
@@
Organization. In Sec. II, we present a compact, implementation-agnostic, physics-compliant system model for DMAs. In Sec. III, we analytically derive upper bounds on the normalized Jacobian and Hessian of the DMA radiation pattern with respect to the DMA configuration, as a function of the DMA’s mutual coupling strength. In Sec. IV, we provide full-wave numerical evidence for the trade-off between strength and simplicity of the mapping from configuration to radiation pattern, and we demonstrate how a stronger sensitivity improves the ability to optimize the channel gain with a DMA. In Sec. V, we use a DMA prototype to experimentally confirm these trends. We close in Sec. VI with a conclusion and a brief outlook to future work.
@@
Notation. $\mathbf{A}=\mathrm{diag}(\mathbf{a})$ denotes the diagonal matrix $\mathbf{A}$ whose diagonal entries are $\mathbf{a}$. $\mathbf{b}=\mathrm{diag}(\mathbf{B})$ denotes the vector $\mathbf{b}$ containing the diagonal entries of the matrix $\mathbf{B}$. $\left[\mathbf{A}^{-1}\right]_{\mathcal{BC}}$ denotes the block of $\mathbf{A}^{-1}$ selected by the sets of indices $\mathcal{B}$ and $\mathcal{C}$. $\left[\mathbf{a}\right]_{\mathcal{B}}$ denotes the vector formed by the elements of $\mathbf{a}$ selected by $\mathcal{B}$. $\mathbf{a}^{\top}$ denotes the transpose of $\mathbf{a}$. $\mathbf{I}_{a}$ denotes the $a\times a$ identity matrix. $\mathbf{0}_{a}$ denotes the $a\times a$ zero matrix.
@@ II System Model
On a high abstraction level, any DMA can be described as consisting of $N_{\mathrm{F}}$ feeds and $N_{\mathrm{M}}$ meta-atoms whose electromagnetic coupling is mediated by waveguides or cavities in a way that is implementation-specific (see Fig. 1). During transmission, the input signal excites the feeds and ultimately the meta-atoms because they are coupled to the feeds; then, the meta-atoms radiate the signal to the scene. During reception, the reverse happens. In this section, we seek to distill the essential wave physics without getting into the weeds of deriving a closed-form expression for analytically evaluating this coupling in a given DMA hardware. We hence do not make any assumption about the DMA hardware specifics other than linearity, passivity, time-invariance and reciprocity, as well as that the meta-atoms are point-like (i.e., very small compared to the operating wavelength). Throughout this paper, we do not explicitly print frequency dependence because we study DMAs in single-frequency operation.
@@ II-A Wave Propagation in the DMA
Our starting point is a description of the $N=N_{\mathrm{F}}+N_{\mathrm{M}}$ primary entities (feeds and meta-atoms) as dipoles. Specifically, one dipole is assigned to the location of each primary entity. The $i$th dipole is a point-like polarizable particle with a specific orientation and of infinitesimal size $\delta_{i}$ that accumulates a charge separation of $q_{i}$ and $-q_{i}$ across its ends when an electromagnetic field $e_{i}$ is applied to it along its orientation. This charge separation gives rise to a dipole moment $p_{i}=q_{i}\delta_{i}$. The tendency of a dipole to build up a dipole moment under the application of an electromagnetic field is characterized by its polarizability: $\alpha_{i}=p_{i}/e_{i}$. The polarizability of the feeds is fixed whereas the polarizability of the meta-atoms is tunable. We define the DMA configuration as
@@
where $\alpha_{0}$ is a reference polarizability value; $\mathcal{M}$ [$\mathcal{F}$] denotes the set of dipole indices associated with meta-atoms [feeds]. Dipoles that have accumulated dipole moments reradiate electromagnetic fields at the same frequency as the initial impinging field. The field at the location and along the orientation of the $j$th dipole originating from a unit dipole moment at the $i$th dipole is defined as the background Green’s function $G_{ji}$. The term “background” highlights that $G_{ji}$ is generally not a free-space Green’s function but specific to the DMA hardware involving waveguides or cavities. The field $e_{i}$ at the location and along the orientation of the $i$th dipole is hence the superposition of the initial impinging field and the reradiated secondary fields:
$e_{i}=\alpha_{i}p_{i}=e_{i}^{\mathrm{inc}}+\sum_{j=1}^{N}G_{ij}p_{j}.$ (2)
@@ II-B DMA Radiation Pattern
Having determined the meta-atoms’ dipole moments via Eq. (3), we now derive the DMA’s radiation pattern. The radiated field has three polarization components ($x$, $y$ and $z$ in Cartesian coordinates); the $k$-component (where $k\in\{x,y,z\}$) of the radiated field at some position $\mathbf{r}$ in the scene is the superposition of the $k$-components of the fields radiated by the meta-atoms to that position: [^5] Implicit in Eq. (7) is that the meta-atoms radiate toward the scene like dipoles in free space (which is a reasonable assumption except for grazing angles) and that there is no static direct contribution from the feeds to the field in the scene (which is reasonable because the feeds are embedded within the DMA rather than patterned onto its front surface like the meta-atoms).
$\begin{split}e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})&=\sum_{i=1}^{N_{\mathrm{M}}}G_{k}^{\mathrm{FS}}(\mathbf{r},\mathbf{r}_{\mathcal{M}_{i}})p_{\mathcal{M}_{i}}(\mathbf{c})=\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\mathbf{p}_{\mathcal{M}}(\mathbf{c})\\ &=\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\left[\left(\mathbf{W}(\mathbf{c})\right)^{-1}\mathbf{e}^{\mathrm{inc}}\right]_{\mathcal{M}}\\ &=\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\left[\left(\mathbf{W}(\mathbf{c})\right)^{-1}\right]_{\mathcal{MF}}\mathbf{e}^{\mathrm{inc}}_{\mathcal{F}},\end{split}$ (7)
where $\mathbf{r}_{\mathcal{M}_{i}}$ is the location of the $i$th meta-atom and $G_{k}^{\mathrm{FS}}(\mathbf{r}_{\mathcal{M}_{i}},\mathbf{r})$ is the projection of the free-space dyadic Green’s tensor between positions $\mathbf{r}_{\mathcal{M}_{i}}$ and $\mathbf{r}$ onto the field component $k$ at position $\mathbf{r}$ and the field component parallel to the orientation of the $i$th dipole at position $\mathbf{r}_{\mathcal{M}_{i}}$ (see Eq. (9.18) in [50] for the closed-form expression), and $\mathbf{v}_{k}(\mathbf{r})=\left[G_{k}^{\mathrm{FS}}(\mathbf{r},\mathbf{r}_{i})\ |\ i\in\mathcal{M}\right]\in\mathbb{C}^{N_{\mathrm{M}}}$. The radiation pattern is hence a linear projection of the meta-atoms’ dipole moments to the scene.
@@
II-C Insights into Mutual Coupling

### v3

@@ Abstract
Dynamic metasurface antennas (DMAs) are a promising embodiment of next-generation reconfigurable antenna technology to realize base stations and access points with reduced cost and power consumption. A DMA is a thin structure patterned on its front with reconfigurable radiating metamaterial elements (meta-atoms) that are excited by waveguides or cavities. Mutual coupling (MC) between the meta-atoms can result in a strongly non-linear dependence of the DMA’s radiation pattern on the configuration of its meta-atoms. However, besides the obvious algorithmic challenges of working with physics-compliant DMA models, it remains unclear how MC in DMAs influences the ability to achieve a desired wireless functionality. In this paper, we provide theoretical, numerical and experimental evidence that strong MC in DMAs increases the radiation pattern sensitivity to the DMA configuration and thereby boosts the available control over the radiation pattern, improving the ability to tailor the radiation pattern to the requirements of a desired wireless functionality. Counterintuitively, we hence encourage next-generation DMA implementations to enhance (rather than suppress) MC, in combination with suitable physics-compliant modeling and optimization. We expect the unveiled mechanism by which MC boosts the radiation pattern control to also apply to other reconfigurable antenna systems based on tunable lumped elements.
@@ I Introduction
However, the scalability of such multi-antenna systems is fundamentally limited by considerations of cost, power consumption, weight, and footprint. To overcome these limitations, next-generation reconfigurable antenna technology seeks to exploit additional degrees of freedom in the physical layer. The software-controllable electromagnetic properties of the resulting hybrid analog/digital architectures enable the dynamic reconfiguration of radiation patterns. Practical implementations are typically based on tunable lumped elements offering electrical tunability at spatially discrete locations. Early examples are electronically steerable passive array radiator (ESPAR) antennas [3, 4, 5, 6, 7, 8, 9], where a few driven antennas are surrounded by tunably loaded parasitic elements. Another important example are reconfigurable pixel-based antennas; these are grids of metallic patches interconnected by tunable lumped elements [10, 11, 12], where some of the patches are the main radiators and the tunable lumped elements control the coupling between all patches. A third example are dynamic metasurface antennas (DMAs); these are surfaces patterned with radiating reconfigurable meta-atoms that are excited by waveguides or cavities [13, 14, 15, 16, 17, 18, 19, 20]. In DMAs, the tunable meta-atoms are simultaneously the tuning mechanism and the radiation mechanism. We focus for concreteness on DMAs in this paper; however, we expect the fundamental principles that we unveil to apply more broadly to any reconfigurable antenna based on tunable lumped elements.
@@
It is relatively well understood that mutual coupling (MC) between the reconfigurable meta-atoms [^1] The question of how MC between physically tunable elements impacts the influence of these physical-layer degrees-of-freedom on system performance is clearly different from the well-studied impact of MC in static arrays on the performance of systems for which only the coherent excitation signals are adjustable [21, 22]. profoundly influences the mathematical nature of the mapping from DMA configuration to DMA radiation pattern. Specifically, this mapping is generally non-linear due to MC, as seen in physics-compliant models for various DMA hardware designs [23, 24, 25, 26]. Because the mapping’s non-linearity severely complicates the algorithmic design, MC is generally considered a vexing nuance. Wireless practitioners hence either avoid these complications by neglecting or deliberately minimizing MC in DMAs to achieve an effectively affine mapping [27, 14, 28, 29, 18, 17, 30, 31, 19, 32], or they develop advanced algorithms based on physics-compliant models that capture MC in DMAs [16, 33, 34]. However, the influence of MC on the strength of the dependence of the DMA radiation pattern on the DMA configuration is generally not considered.
Recently, inspired by earlier work on the benefits of multi-path propagation in wireless localization [35], we presented preliminary full-wave numerical evidence that MC boosts the influence of the DMA configuration on the DMA radiation pattern [36]; we further suggested that stronger MC should therefore improve the ability to tailor the DMA radiation pattern to a desired wireless functionality [36]. This generally overlooked effect of MC in DMAs can drastically change the assessment of the role of MC in DMAs. In essence, a trade-off between strength and simplicity of the dependence of the DMA radiation pattern on the DMA configuration emerges. Intuitively, these two effects of MC can be understood in terms of multi-bounce paths between the meta-atoms [^2] The relation between physics-compliant models and a multi-bounce-path picture has been worked out in the context of RIS-parametrized radio environments using a coupled-dipole formalism in [37] and using multi-port network theory in Sec. II.A of [38].: The stronger the MC in the DMA is, the more the wave bounces back and forth between meta-atoms, resulting in an accumulation of more sensitivity to the DMA configuration, which in turn yields a i) stronger and ii) more strongly non-linear (i.e., less simple) dependence of the DMA radiation pattern on the DMA configuration. Consequently, it may be desirable to deploy DMA hardware with enhanced (rather than suppressed) MC in combination with suitable physics-compliant models and optimization techniques.
@@
    MC strength between the DMA’s meta-atoms,
@@
  • We formulate a compact, generic (implementation-agnostic), physics-compliant DMA model within the coupled-dipole framework. Our formulation collapses architecture specifics into a background coupling operator, so that our formulation is architecture-agnostic and maximally compact. [^3] A related, similarly compact formulation in terms of impedance parameters was recently proposed in [39]; the first preprint of [39] appeared shortly after the first preprint of our present manuscript. Our formulation transparently exposes the mathematical role of MC in DMAs. We also clarify how to recover common simplified linear models upon neglecting MC. We further expect our formulation’s compactness and generality to be ideal for future efforts to experimentally estimate the model parameters (see discussion in Sec. VI). Our formulation complements (rather than supersedes) existing architecture-specific coupled-dipole formulations like [24, 25] which remain indispensable to construct the background coupling operator in numerical studies.
@@
  • We derive analytical expressions and upper bounds for the normalized Jacobian and Hessian of the DMA radiation pattern with respect to the DMA configuration. The upper bounds monotonically increase with the MC strength.
@@
  • We systematically examine in full-wave simulations how the four pivotal quantities listed above depend on the MC strength in a DMA. We reveal clear trends of positive correlation between MC strength and sensitivity, achieved channel gain enhancement, and non-linearity of the mapping from configuration to radiation pattern.
@@
  • We conduct experiments with a DMA prototype operating in the lower K-band with 1-bit programmable meta-atoms. We introduce an original time-gating technique to adjust the effective MC strength in post-processing. We report clear experimental evidence for the numerically observed trends.
@@
Organization. In Sec. II, we present a compact, implementation-agnostic, physics-compliant system model for DMAs. In Sec. III, we analytically derive upper bounds on the normalized Jacobian and Hessian of the DMA radiation pattern with respect to the DMA configuration, as a function of the DMA’s MC strength. In Sec. IV, we provide full-wave numerical evidence for the trade-off between strength and simplicity of the mapping from configuration to radiation pattern, and we demonstrate how a stronger sensitivity improves the ability to optimize the channel gain with a DMA. In Sec. V, we use a DMA prototype to experimentally confirm these trends. We close in Sec. VII with a conclusion and a brief outlook to future work.
@@
Notation. $\mathbf{A}=\mathrm{diag}(\mathbf{a})$ denotes the diagonal matrix $\mathbf{A}$ whose diagonal entries are $\mathbf{a}$. $\mathbf{b}=\mathrm{diag}(\mathbf{B})$ denotes the vector $\mathbf{b}$ containing the diagonal entries of the matrix $\mathbf{B}$. $\left[\mathbf{A}^{-1}\right]_{\mathcal{BC}}$ denotes the block of $\mathbf{A}^{-1}$ selected by the sets of indices $\mathcal{B}$ and $\mathcal{C}$. $\left[\mathbf{a}\right]_{\mathcal{B}}$ denotes the vector formed by the elements of $\mathbf{a}$ selected by $\mathcal{B}$. $\mathbf{a}^{\top}$ denotes the transpose of $\mathbf{a}$. $\mathbf{I}_{a}$ denotes the $a\times a$ identity matrix. $\mathbf{0}_{a}$ denotes the $a\times a$ zero matrix. $\mathbf{0}_{\mathcal{BC}}$ denotes the $|\mathcal{B}|\times|\mathcal{C}|$ zero matrix, with $|\mathcal{B}|$ denoting the cardinality of $\mathcal{B}$.
@@ II System Model
At a high level of abstraction, any DMA can be described as consisting of $N_{\mathrm{F}}$ feeds and $N_{\mathrm{M}}$ meta-atoms whose electromagnetic coupling is mediated by waveguides or cavities in a way that is implementation-specific (see Fig. 1). During transmission, the input signal excites the feeds and ultimately the meta-atoms because they are coupled to the feeds; then, the meta-atoms radiate the signal to the scene. During reception, the reverse happens.
Existing physics-compliant DMA models (e.g., [40, 23, 24, 25, 26]) seek to explicitly describe each detail of a given DMA architecture in closed form to determine the coupling between feeds and meta-atoms. Thereby, they enable numerical studies such as ours in Sec. IV. However, thereby they are also inevitably specific to a given DMA architecture, their experimental applicability is vulnerable to fabrication inaccuracies, and high-level conceptual insights are somewhat obscured (such as the fundamental trade-off between strength and simplicity of the dependence of the radiation pattern on the configuration that we investigate here).
In this section, we seek to distill the essential wave physics without getting into the weeds of deriving a closed-form expression for analytically evaluating this coupling in a given DMA hardware. We hence do not make any assumption about the DMA hardware specifics such that our compact model applies to both types of common DMA implementations shown in Fig. 1. We only assume that the DMA’s static parts are linear, passive, time-invariant and reciprocal, which is generally the case, and that the DMA’s tunable parts (i.e., the meta-atoms) are very small compared to the operating wavelength, which is justified by their intrinsic sub-wavelength nature and an accurate approximation according to earlier work [40, 23, 24, 25]. Throughout this paper, we do not explicitly print frequency dependence because we study DMAs in single-frequency operation.
@@ II-A Wave Propagation in the DMA
Our starting point is a description of the $N=N_{\mathrm{F}}+N_{\mathrm{M}}$ primary entities (feeds and meta-atoms) as dipoles. The validity of this approach was validated in commercial simulation software and experiments in [40, 23, 24, 25]. Specifically, one dipole is assigned to the location of each primary entity. The $i$th dipole is a point-like polarizable particle with a specific orientation and of infinitesimal size $\delta_{i}$ that accumulates a charge separation of $q_{i}$ and $q_{i}$ across its ends when an electromagnetic field $e_{i}$ is applied to it along its orientation. This charge separation gives rise to a dipole moment $p_{i}=q_{i}\delta_{i}$. The tendency of a dipole to build up a dipole moment under the application of an electromagnetic field is characterized by its polarizability: $\alpha_{i}=p_{i}/e_{i}$. The polarizability of the feeds is fixed whereas the polarizability of the meta-atoms is tunable. We define the DMA configuration as
@@
where $\alpha_{0}$ is a reference polarizability value; $\mathcal{M}$ $\mathcal{F}$] denotes the set of dipole indices associated with meta-atoms [feeds]. Dipoles that have accumulated dipole moments reradiate electromagnetic fields at the same frequency as the initial impinging field. The field at the location and along the orientation of the $j$th dipole originating from a unit dipole moment at the $i$th dipole is defined as the background Green’s function $G_{ji}$. The term “background” highlights that $G_{ji}$ is generally not a free-space Green’s function but specific to the DMA hardware involving waveguides or cavities. The field $e_{i}$ at the location and along the orientation of the $i$th dipole is hence the superposition of the initial impinging field and the reradiated secondary fields:
$e_{i}=\alpha_{i}^{-1}p_{i}=e_{i}^{\mathrm{inc}}+\sum_{j=1}^{N}G_{ij}p_{j}.$ (2)
@@ II-B DMA Radiation Pattern
Having determined the meta-atoms’ dipole moments via Eq. (3), we now derive the DMA’s radiation pattern. As explained in [25] (see discussion around (21) and (22) therein) as well as [52], the radiated field can be evaluated by treating the meta-atoms as dipoles in free space, upon scaling the dipole moments found in (3) by a real-valued scalar $\nu$. This scaling accounts, among other effects, for the electromagnetic images of the dipoles. For the analysis in this paper, the value of this scaling factor is immaterial. The radiated field has three polarization components ($x$, $y$ and $z$ in Cartesian coordinates); the $k$-component (where $k\in\{x,y,z\}$) of the radiated field at some position $\mathbf{r}$ in the scene is the superposition of the $k$-components of the fields radiated by the meta-atoms to that position: [^6] Implicit in Eq. (7) is that the meta-atoms radiate toward the scene like dipoles in free space (which is a reasonable assumption except for grazing angles) and that there is no static direct contribution from the feeds to the field in the scene (which is reasonable because the feeds are embedded within the DMA rather than patterned onto its front surface like the meta-atoms).
$\begin{split}e^{\mathrm{rad},k}(\mathbf{r},\mathbf{c})&=\nu\sum_{i=1}^{N_{\mathrm{M}}}G_{k}^{\mathrm{FS}}(\mathbf{r},\mathbf{r}_{\mathcal{M}_{i}})p_{\mathcal{M}_{i}}(\mathbf{c})=\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\mathbf{p}_{\mathcal{M}}(\mathbf{c})\\ &=\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\left[\left(\mathbf{W}(\mathbf{c})\right)^{-1}\mathbf{e}^{\mathrm{inc}}\right]_{\mathcal{M}}\\ &=\left(\mathbf{v}_{k}(\mathbf{r})\right)^{\top}\left[\left(\mathbf{W}(\mathbf{c})\right)^{-1}\right]_{\mathcal{MF}}\mathbf{e}^{\mathrm{inc}}_{\mathcal{F}},\end{split}$ (7)
where $\mathbf{r}_{\mathcal{M}_{i}}$ is the location of the $i$th meta-atom and $G_{k}^{\mathrm{FS}}(\mathbf{r},\mathbf{r}_{\mathcal{M}_{i}})$ is the projection of the free-space dyadic Green’s tensor between positions $\mathbf{r}_{\mathcal{M}_{i}}$ and $\mathbf{r}$ onto the field component $k$ at position $\mathbf{r}$ and the field component parallel to the orientation of the $i$th dipole at position $\mathbf{r}_{\mathcal{M}_{i}}$ (see Eq. (9.18) in [53] for the closed-form expression), and $\mathbf{v}_{k}(\mathbf{r})=\nu\left[G_{k}^{\mathrm{FS}}(\mathbf{r},\mathbf{r}_{i})\ |\ i\in\mathcal{M}\right]\in\mathbb{C}^{N_{\mathrm{M}}}$. The radiation pattern is hence a linear projection of the meta-atoms’ dipole moments to the scene.
@@
II-C Insights into MC